# Bayes, E-values, and Testing

### v3

@@ B.1 Uniqueness of the log-score transformation
Among continuous order-reversing transforms, the negative logarithm $x\mapsto-\log x$ is the unique map (up to positive scaling) that converts multiplicative likelihood ratios into additive evidence increments.
@@
B.2 The evidence cone as a convex algebra
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}}_{H_{0}})$ be a filtered probability space under $H_{0}$. We collect the algebraic closure properties of the E-process cone $\mathcal{E}^{\mathrm{proc}}(H_{0})$ stated in Theorem˜1.1 and prove them in a self-contained manner.
@@ B.2 The evidence cone as a convex algebra
These results establish that the class of E-processes forms a convex cone closed under predictable stopping and countable convex mixtures, as stated in Theorems˜1.1 and 4.2. These convex-algebraic properties are compatible with the fork-convexity geometry studied in the betting and e-process literature [31, Sec. 7.2].
@@ B.4 Computational boundary: an impossibility result
For general e-process constructions beyond product-of-stepwise e-values, see Ramdas [31, Ch. 7], including the distinction between test supermartingales and e-processes and several general constructions for composite testing problems.
@@
The distinction between test (super)martingales (including product-form constructions) and the more general notion of e-processes is emphasized in Ramdas [31, Def. 7.3], which also records the nontrivial equivalence between stopping-time formulations and domination by test martingale families (proved via Snell envelopes and Doob decompositions in Ramdas et al. 32). See Ramdas [31, Sec. 7.2–7.4 and bibliographical note] for historical context and for examples (e.g. exchangeability testing) where e-processes exhibit behavior not captured by test supermartingales; geometric closure notions such as fork-convexity play a key role. Our convex-algebraic closure properties (Section˜B.2) are compatible with, and complementary to, the fork-convexity geometry emphasized in the betting literature.
@@ C.4 Simulation: evidence growth under correct and misspecified alternatives
Our LR representation results are representation statements (what evidence must look like under coherence / Bayes-risk structure), not claims that all useful e-processes must be built as products of one-step e-values. General e-process constructions—including those not representable as multiplicative test martingales for certain composite or nonparametric problems—are discussed in the game-theoretic safe-testing literature; see, e.g., Ramdas et al. [33]. The point here is orthogonal: whenever a coherent predictive state exists, the induced evidence process is an LR (or mixture of LRs), and boundary choice is governed by Bayes risk rather than Markov calibration.

### v4

@@ B.1 Uniqueness of the log-score transformation
Among continuous order-preserving transforms, the negative logarithm $x\mapsto-\log x$ is the unique map (up to positive scaling) that converts multiplicative likelihood ratios into additive evidence increments.
@@
B.2 The evidence class as a convex algebra
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}}_{H_{0}})$ be a filtered probability space under $H_{0}$. We collect the algebraic closure properties of the E-process class $\mathcal{E}^{\mathrm{proc}}(H_{0})$ stated in Theorem˜1.1 and prove them in a self-contained manner.
@@ B.2 The evidence class as a convex algebra
These results establish that the class of E-processes forms a convex set, closed under scaling by $c\in(0,1]$, predictable stopping, and countable convex mixtures, as stated in Theorems˜1.1 and 4.3. Note that this class is not a cone in the algebraic sense: multiplication by constants greater than one generally violates the unit-expectation constraint $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$, so only scaling by $c\in(0,1]$ preserves membership. These convex-algebraic properties are compatible with the fork-convexity geometry studied in the betting and e-process literature [33, Sec. 7.2].
Proposition B.7 (Layer separation).
The following are logically distinct objects:
  1. 1.
    A likelihood-ratio representation $E=dQ/dP$ (representation layer);
  2. 2.
    A supermartingale property $\operatorname{\mathbb{E}}_{P}[E_{t}\mid\mathcal{F}_{t-1}]\leq E_{t-1}$ (validity layer);
  3. 3.
    A boundary rule $\tau=\inf\{t:E_{t}\geq b\}$ (decision layer).
Each may be specified independently of the others, and optimality statements at one layer do not imply optimality at another. In particular, a process may possess a likelihood-ratio representation without being a supermartingale (if the denominator is misspecified), and a valid e-process may admit no likelihood-ratio representation at all.
Proof.
Independence of (1) and (2): any ratio $dQ/dP$ is a $P$-martingale when $Q\ll P$, but ratios formed under a misspecified model (i.e., $Q$ not absolutely continuous on the filtration) need not be supermartingales. Conversely, e-processes constructed via convex mixtures (Propositions˜B.4 and B.6) need not admit a single likelihood-ratio representation. Independence of (3): the boundary $b$ is a free parameter that controls the type-I/power trade-off but does not affect validity or the representational structure of $(E_{t})$. ∎
Remark B.8 (Nonparametric validity).
Uniform validity over a null class $\mathcal{P}$ is a statement about the certificate (validity) layer: $\sup_{P\in\mathcal{P}}\operatorname{\mathbb{E}}_{P}[E_{t}\mid\mathcal{F}_{t-1}]\leq E_{t-1}$. It does not constrain the alternative-side geometry of $\log E_{t}$, which may inherit KL/MDP structure when a likelihood-ratio representation exists (Proposition˜B.9).
@@ B.4 Computational boundary: an impossibility result
For general e-process constructions beyond product-of-stepwise e-values, see Ramdas [33, Ch. 7], including the distinction between test supermartingales and e-processes and several general constructions for composite testing problems.
@@
The distinction between test (super)martingales (including product-form constructions) and the more general notion of e-processes is emphasized in Ramdas [33, Def. 7.3], which also records the nontrivial equivalence between stopping-time formulations and domination by test martingale families (proved via Snell envelopes and Doob decompositions in Ramdas et al. 34). See Ramdas [33, Sec. 7.2–7.4 and bibliographical note] for historical context and for examples (e.g. exchangeability testing) where e-processes exhibit behavior not captured by test supermartingales; geometric closure notions such as fork-convexity play a key role. Our convex-algebraic closure properties (Section˜B.2) are compatible with, and complementary to, the fork-convexity geometry emphasized in the betting literature.
@@ C.4 Simulation: evidence growth under correct and misspecified alternatives
Our LR representation results are representation statements (what evidence must look like under coherence / Bayes-risk structure), not claims that all useful e-processes must be built as products of one-step e-values. General e-process constructions—including those not representable as multiplicative test martingales for certain composite or nonparametric problems—are discussed in the game-theoretic safe-testing literature; see, e.g., Ram [1]. The point here is orthogonal: whenever a coherent predictive state exists, the induced evidence process is an LR (or mixture of LRs), and boundary choice is governed by Bayes risk rather than Markov calibration.